# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ 4.2 Continuation of the local solution
Theorem 4.1 provides the explicit local solution to the slow system (12) on the interval $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$. When $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$, what can we say about the slow solution and its time derivative at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$? The next result shows that the slow solution is continuous at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ but its time derivative is not.
@@
See Appendix B.2 for the proof. ∎
4.3 Explicit global solution
So far, we have computed the explicit local solution of the slow system (12) on an interval $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ which, when $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is finite, can be extended to $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. We can apply Theorem 4.1 and Lemma 4.1 again with the initial condition $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ to extend the local solution to $[0,\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$. This is because $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ and all assumptions of Theorem 4.1 hold.
@@ 4.3 Explicit global solution
$\boldsymbol{p}^{(k)}$ $=\boldsymbol{p}^{(k-1)}+\begin{cases}\min\left(\Delta^{(k-1)},1/t\right)\boldsymbol{d}^{(k-1)}&\text{if}\ t>0\\ \Delta^{(k-1)}\boldsymbol{d}^{(k-1)}&\text{if $t=0$ and $\Delta^{(k-1)}<+\infty$},\\ \boldsymbol{0}&\text{otherwise.}\end{cases}$
The slow solution is given piecewise on the intervals $[\tau^{(k-1)},\tau^{(k)})$:
@@
This gives the explicit solution of the slow system over $[0,\tau^{(k)})$ up to any $k\in\mathbb{N}$. Note that if $t>0$, then the minimum in (29) is attained at $1/t\iff\tau^{(K)}=+\infty$ for some $K\in\mathbb{N}$.
What can be said about the asymptotic limit $k\to+\infty$? The following Theorem asserts the limit converges in finite time, in the sense that there exists some nonnegative integer $K$ such that $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ over the interval $[\tau^{(K)},+\infty)$.
@@
Let $t\geqslant 0$ and $\boldsymbol{p}^{(0)}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. Consider the slow system (12) with initial condition $\boldsymbol{p}(0)=\boldsymbol{p}^{(0)}$, whose solution is given by (30). Then there exists a nonnegative integer $K$ such that $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ and $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ on $\tau\in[\tau^{(K)},+\infty)$.
@@
We will show there exists a nonnegative integer $K$ for which $\boldsymbol{d}^{(K)}=\boldsymbol{0}$. This suffices because the projection onto the closed, convex set $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ is unique (see A.17), and so the subgradients are different if and only if their norms are different.
@@
Let $k$ be a positive integer and suppose $\tau^{(j)}<+\infty$ for every $j\in\{1,\dots,k\}$ in (29). On the one hand, Lemma 4.1 implies $\left\|{\boldsymbol{d}^{(j)}}\right\|_{2}<\left\|{\boldsymbol{d}^{(j-1)}}\right\|_{2}$. Hence the descent directions $\{\boldsymbol{d}^{(1)},\dots,\boldsymbol{d}^{(k)}\}$ are obtained from projections landing all on different faces of the convex cone $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. On the other hand, since $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ has at most $4^{n}$ faces and $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})$, there is at least one face on which the projection is zero. Let $K$ denote the smallest such integer in the construction (30). Then $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ and $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ on $[\tau^{(K)},+\infty)$. ∎
Remark 4.1 (Drift to infinity, again).
@@
Without the full row rank assumption $\mathrm{rank}({\bm{A}})=m$ (more generally, $\boldsymbol{b}\in\text{Im}(\bm{A})$), there may be some $K\in\mathbb{N}$ for which $\boldsymbol{d}^{(K-1)}\neq 0$ with $\bm{A}^{\top}\boldsymbol{d}^{(K-1)}=\boldsymbol{0}$, so that $\Delta^{(K-1)}=+\infty$ and $\tau^{(K)}=+\infty$. This does not cause an issue when $t>0$ since $\boldsymbol{p}^{(K)}$ remains finite. However, when $t=0$, we have $\lim_{\tau\to+\infty}\left\|{\boldsymbol{p}(\tau)}\right\|_{2}=+\infty$, which means the trajectory of the slow system diverges and there are no feasible solutions to the corresponding (BP) and (dBP) problems.
4.4 Recovery of solutions to basis pursuit denoising
@@
The analysis in Sections 4.14.3 yields the global solution to the slow system (12). Crucially, its asymptotic limit can be computed explicitly from (30) because it converges in finite time, meaning $\lim_{\tau\to+\infty}\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ for some nonnegative integer $K$. This recovers a pair of primal and dual solutions $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ to BPDN. Indeed, the minimal selection princple implies $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{p}^{(K)}$, while a primal solution follows from the optimality conditions (6) and Lemma 3.1: Letting $\hat{\boldsymbol{u}}^{(K)}$ denote an optimal solution to the NNLS problem (16), then
$-t\boldsymbol{p}^{(K)}=\boldsymbol{b}-\bm{A}\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}\implies\boldsymbol{x}^{s}(t,\boldsymbol{b})=\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}.$
4.5 An active set algorithm for basis pursuit denoising
@@
Taken together, our results yield Algorithm 1, presented in the introduction of this work. It computes the trajectory of slow system (12) until the asymptotic limit is reached in finite time. Thanks to (29) and (30), the trajectory is integrated exactly because the descent direction and timestep can be computed using, e.g., active set NNLS or cone projection algorithms as per Remark 3.2 in Section 3.2.
@@
Algorithm 1 requires some input $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ with $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}=1$ to serve as the initial condition of the slow system (12). For this, one can take $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. In addition, Algorithm 1 is naturally suited for computing regularization paths. To do so, one selects a decreasing sequence of nonnegative hyperparameters $\{t^{(l)}\}_{l=0}^{L}$ and sequentially computes the dual solutions $\{\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})\}_{l=0}^{L}$, using $\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})$ as the input in Algorithm 1 for computing $\boldsymbol{p}^{s}(t^{(l+1)},\boldsymbol{b})$. Again, one can start with $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$ and $\boldsymbol{p}^{s}(t^{(0)},\boldsymbol{b})=-\boldsymbol{b}/t^{(0)}$.
5 Perturbation analysis of the slow system
Sections 34 describes how BPDN is characterized from the dynamics of the slow system (12) and its slow solution (30). In particular, the analysis yields an exact algorithm that computes a pair of primal and dual solutions to BPDN for fixed hyperparameter $t$ and data $\boldsymbol{b}$. What happens, however, to the dynamics of the slow system (12) when the hyperparameter or data or both are perturbed? This matter in practice if, for example, one seeks to assess robustness of solutions with respect to the hyperparameter and data [21, 46, 28]. Another example is if the desired hyperparameter is unknown. In this case, one typically compute a regularization path or the entire solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ and then select the desired hyperparameter, e.g., based on the sparsity of the solution or data-driven methods [30, 50, 25, 22, 11]. In any case, the perturbation analysis of BPDN is crucial to tackle these questions. In the context of minimal selection principle, this amounts to analyzing the dependence on parameters of the slow system and its slow solution.
@@ 5 Perturbation analysis of the slow system
Motivated by these matters, we characterize here how the dynamics of the slow system and its slow solution behave under perturbations on the hyperparameter $t$ and data $\boldsymbol{b}$. First, we present in Section 5.1 how the descent direction of the slow system behaves under perturbations. We then use this in conjunction with the minimal selection principle in Section 5.2 to describe the local dependence on the parameters $t$ and $\boldsymbol{b}$ of the slow solution and, therefore, the primal and dual solutions to BPDN. As we argue in Section 5.3, this local dependence leads to a possibly non-unique local continuation of the primal solution to BPDN. This non-uniqueness issue, a well-known problem in the literature [63], arises from the non-uniqueness of solutions to an NNLS problem, as also previously reported by [11]. Finally, we present in Section 5.4 the global dependence on parameters of the slow solution, naturally yielding a rigorous homotopy algorithm, similar to the one proposed in [11] but based firmly on the minimal selection principle, for computing the solution path to BPDN.
5.1 Perturbations of the descent direction
@@
We first describe how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ changes under perturbations in the hyperparameter and data. This will be used in Section 5.2 to characterize the local dependence on parameters of the slow solution (30).
@@
Let $t_{0}\geqslant 0$, let $\boldsymbol{p}_{0}\in\mathrm{dom~{}}V(\cdot;t_{0},\boldsymbol{b})$, and let $\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ be a global minimum of the NNLS problem (16) so that $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-(\boldsymbol{b}+t_{0}\boldsymbol{p}_{0})$. In addition, let $\delta_{0}\in\mathbb{R}$ be such that $t_{0}+\delta_{0}\geqslant 0$ and let $\boldsymbol{\delta}\boldsymbol{b}\in\mathbb{R}^{m}$.
@@
    The perturbed descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})$ is given by
    $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})=\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t,\boldsymbol{b},\delta_{0},\boldsymbol{\delta}\boldsymbol{b})-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0},$ (31)
@@
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b},\delta_{0},\boldsymbol{\delta}\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$ (32)
    subject $\qquad\qquad\text{subject to}\ \begin{cases}\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}_{0})}&\geqslant-\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b}),\\ \boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}&=\boldsymbol{0}.\end{cases}$
@@
    (Linear perturbations in hyperparameter) Suppose $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$, let $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$, and let $\delta_{0}\boldsymbol{b}=-(t_{0}-t)\boldsymbol{q}$ with $\boldsymbol{q}\in\mathbb{R}^{m}$. Then (31) and (32) refine to
    $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=(1-t/t_{0})\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}_{0}+\boldsymbol{q})\right)$ (33)
@@
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}+t_{0}(\boldsymbol{p}_{0}+\boldsymbol{q})}\right\|_{2}^{2}$ (34)
    subject $\text{subject to}\ \begin{cases}&\boldsymbol{v}_{j}\geqslant-\hat{\boldsymbol{u}}_{j}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})/(1-t/t_{0})\ \text{if $j\in\mathcal{E}(\boldsymbol{p}_{0})$},\\ &\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}.\end{cases}$
@@
5.2 Local dependence on parameters of the slow solution
In the stationary case $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$ with $t_{0}>0$, Lemma 5.1 characterizes implicitly how the primal and dual solutions to BPDN change under perturbations in the hyperparameter and data. Here, we show how these perturbations characterize these changes explicitly.

### v2

@@ 4.3 Continuation of the local solution
Theorem 4.1 provides the explicit local solution of the slow system (12) on the interval $0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$. When $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$, what can we say about the slow solution and system at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$?
@@
Hence we can extend the analysis done in Theorem 4.1 to the value $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$, yielding $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))=\boldsymbol{p}_{a}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$. Thus (23) holds at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$.
Next, we prove inequality (28). Note that Proposition 4.1(iii) immediately implies
$\left\|{\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})}\right\|_{2}\leqslant e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}.$
In addition, since the projection onto the closed and convex set $\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})$ is unique (A.17), it sufficies to show that $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Now, apply Theorem 4.1 with the initial value $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ to find
on the interval $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$. Suppose, for a contradiction, that
$\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})=e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Then (29) can be developed into
However, this implies $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})+f(\tau-\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}>\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ for every $\tau\in(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$, contradicting the definition of $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ in (18). We therefore conclude $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. ∎
4.4 Explicit global solution of the slow system
So far, we have computed the explicit local solution of the slow system (12) on an interval $0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, which can be extended to $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ when the value is finite. We can apply Theorem 4.1 and Lemma 4.1 again with the initial condition $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ to extend the local solution to $0,\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b}))$. This is because $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ and all assumptions of Theorem 4.1 hold.
@@ 4.4 Explicit global solution of the slow system
[…]
The slow solution is given piecewise on the intervals $\tau^{(k-1)},\tau^{(k)})$:
@@
This gives the explicit solution of the slow system over $0,\tau^{(k)})$ up to any $k\in\mathbb{N}$. Note that if $t>0$, then the minimum in (30) is attained at $1/t\iff\tau^{(K)}=+\infty$ for some $K\in\mathbb{N}$.
What can be said about the asymptotic limit $k\to+\infty$? The following Theorem asserts the limit converges in finite time, in the sense that there exists some nonnegative integer $K$ such that $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ over the interval $\tau^{(K)},+\infty)$.
@@
Let $t\geqslant 0$ and $\boldsymbol{p}^{(0)}\in\mathrm{dom~}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. Consider the slow system (12) with initial condition $\boldsymbol{p}(0)=\boldsymbol{p}^{(0)}$, whose solution is given by (31). Then there exists a nonnegative integer $K$ such that on the interval $\tau\in[\tau^{(K)},+\infty)$,
$\boldsymbol{p}(\tau)=\begin{cases}\boldsymbol{p}^{(K)}+f(t,\tau-\tau^{(K)})\boldsymbol{d}^{(K)}&\,\text{if $t>0$},\\ \boldsymbol{p}^{(K)}&\,\text{if $t=0$}.\end{cases}$
@@
We will show there exists a nonnegative integer $K$ such that $t\Delta^{(K)}\geqslant 1$ when $t>0$ or $\bm{A}^{\top}\boldsymbol{d}^{(K)}=\boldsymbol{0}$ when $t=0$. Let $k$ be a positive integer and suppose $\tau^{(j)}<+\infty$ for every $j\in\{1,\dots,k\}$ in (30). First, note that the directions $\{\boldsymbol{d}^{(1)},\dots,\boldsymbol{d}^{(k)}\}$ are obtained from projections landing on different faces of the convex cone $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. Indeed, this follows because Lemma 4.1 implies $\left\|{\boldsymbol{d}^{(j)}}\right\|_{2}<\left\|{\boldsymbol{d}^{(j-1)}}\right\|_{2}$ for $j\in\{1,\dots,k\}$ and because the projection onto the closed, convex set $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ is unique (see A.17). Since $\mathrm{rank}({\bm{A}})=m$ by assumption, there is at least one face on which the norm of the projection onto $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ is equal to zero. Thus there exists some $K>k$ such that either $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ or, when $t>0$, $t\Delta^{(K)}\geqslant 1$, whichever happens first (recall from Proposition 4.1(i) that $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ if and only if $\Delta^{(K)}=+\infty$). This yields the desired result. ∎
Remark 4.1.
@@
The proof above shows that finite-time convergence holds even when $\boldsymbol{b}\notin\text{Im}(\bm{A})$. However, in that case there may be some $K\in\mathbb{N}$ for which $\boldsymbol{d}^{(K-1)}\neq 0$ with $\bm{A}^{\top}\boldsymbol{d}^{(K-1)}=\boldsymbol{0}$ (this will always be the case when $t=0$). This causes no issues when $t>0$ because $\boldsymbol{p}^{(K)}$ remains finite. However, when $t=0$ we have $\lim_{\tau\to+\infty}\left\|{\boldsymbol{p}(\tau)}\right\|_{2}=+\infty$, which means the slow solution diverges and there are no feasible solutions to the corresponding (BP) and (dBP) problems.
4.5 An exact algorithm for recovering optimal solutions to the lasso problem
@@
The analysis in Sections 4.24.4 yields the global solution to the slow system (12). Crucially, its asymptotic limit can be computed explicitly from (31) because it converges in finite time, meaning $\lim_{\tau\to+\infty}\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ for some nonnegative integer $K$. This recovers a pair of primal and dual lasso solutions $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$. Indeed, the minimal selection principle implies $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{p}^{(K)}$, while a primal solution follows from the optimality conditions (6) and Lemma 3.1: Letting $\hat{\boldsymbol{u}}^{(K)}$ denote an optimal solution to the NNLS problem (16), then
$t\boldsymbol{p}^{(K)}=\boldsymbol{b}-\bm{A}\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}\implies\boldsymbol{x}^{s}(t,\boldsymbol{b})=\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}.$
@@
Our results yield Algorithm 1, presented in the introduction of this paper. The algorithm integrates the slow system (12) and computes its asymptotic limit in a finite number of steps. The slow solution is calculated exactly because the descent directions and timesteps can be computed using, e.g., active set NNLS or cone projection algorithms as per Remark 3.2.
@@
Note that Algorithm 1 requires an input $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ with $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}\leqslant 1$. For this, one can always take $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. In addition, Algorithm 1 is particularly well suited for computing regularization paths. To do so, one selects a decreasing sequence of nonnegative hyperparameters $\{t^{(l)}\}_{l=0}^{L}$ (starting perhaps from $\boldsymbol{p}^{s}(t^{(0)},\boldsymbol{b})=-\boldsymbol{b}/t^{(0)}$) and sequentially computes the dual solutions $\{\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})\}_{l=0}^{L}$, using $\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})$ as the input in Algorithm 1 for computing $\boldsymbol{p}^{s}(t^{(l+1)},\boldsymbol{b})$.
5 Continuation in the hyperparameter of the asymptotic limit
Sections 34 provide a novel characterization of the lasso problem using the dynamics of the slow system (12) and its slow solution (31) for fixed hyperparameter $t$. In practice, one often seeks to compute the solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ to, e.g., assess robustness of solutions [16, 23] or select a hyperparameter using data-driven methods [25, 38, 20, 17, 9]. This historically motivated the development of the LARS algorithm [38, 20], which is now well-known to fail without technical assumptions that are difficult to verify [9, Proposition 4.1].
@@ 5 Continuation in the hyperparameter of the asymptotic limit
Here, we present a rigorous analysis of continuation of the asymptotic limit of the slow system (12) in terms of the hyperparameter $t$, yielding lasso solution paths. We present in Section 5.1 the local dependence of the slow solution on $t$. As we argue in Section 5.2, this local dependence leads to a possibly non-unique local continuation of the primal lasso solution. This non-uniqueness issue, a well-known problem in the literature [48], arises from the non-uniqueness of solutions to an NNLS problem, as previously reported by [9]. Finally, we present in Section 5.3 the global dependence of the slow solution, on $t$, naturally yielding a rigorous homotopy algorithm based on the minimal selection principle.
5.1 Local dependence on hyperparameter of the slow solution
@@
We first describe how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ changes under perturbations in the hyperparameter and data.
@@
Let $t_{0}\geqslant 0$, let $\boldsymbol{p}_{0}\in\mathrm{dom~}V(\cdot;t_{0},\boldsymbol{b})$, and let $\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ be a global minimum of the NNLS problem (16) so that $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-(\boldsymbol{b}+t_{0}\boldsymbol{p}_{0})$. In addition, let $\delta_{0}\in\mathbb{R}$ be such that $t_{0}+\delta_{0}\geqslant 0$. Then:
@@
    The perturbed descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b})$ is given by
    $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b})=\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t,\boldsymbol{b},\delta_{0},\boldsymbol{\delta}\boldsymbol{b})-\delta_{0}\boldsymbol{p}_{0},$ (32)
@@
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b},\delta_{0})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$ (33)
[…]
@@
    (Linear perturbations in hyperparameter) Suppose $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$ and let $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$. Then (32) and (33) refine to
    $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=(1-t/t_{0})\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b})+t_{0}\boldsymbol{p}_{0}\right)$ (34)
@@
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}+t_{0}(\boldsymbol{p}_{0})}\right\|_{2}^{2}$ (35)
[…]
@@
In the stationary case $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$ with $t_{0}>0$, Lemma 5.1 characterizes implicitly how the primal and dual lasso solutions change under perturbations in the hyperparameter. Next, we show how these perturbations characterize these changes explicitly.